# Soft Masked Transformer for Point Cloud Processing with Skip Attention-Based Upsampling

### v1

@@ V Conclusion
In this paper, we introduced a novel Soft Masked Transformer to effectively capture contextual and task-specific information from point clouds. Additionally, we proposed a Skip Attention-based up-sampling block to integrate features from different resolution points across encoding layers. Furthermore, we presented a Shared Position Encoding strategy. By incorporating these modules, we constructed an SMTransformer network. Our method was evaluated across various tasks, including indoor and outdoor semantic segmentation and classification. Through extensive experiments on challenging benchmarks, thorough ablation studies and theoretical analysis, we demonstrated the robustness and effectiveness of our approach on real-world datasets. Our contributions significantly advance the state-of-the-art in-point cloud processing. The introduced techniques, including the Soft Masked Transformer, Skip Attention-based Up-sampling block, and Shared Position Encoding strategy, provide notable improvements in capturing intricate details and enhancing the performance of point cloud processing tasks. Experimental results on diverse datasets confirmed the efficacy of our proposed approach. As future work, exploring additional priors or refining the network architecture could offer promising avenues to further improve point cloud processing.

### v2

@@ V Conclusion
In this paper, we introduced a novel Soft Masked Transformer to effectively capture contextual and task-specific information from point clouds. Additionally, we proposed a Skip Attention-based up-sampling block to integrate features from different resolution points across encoding layers. Furthermore, we presented a Shared Point Position Encoding strategy. By incorporating these modules, we constructed an SMTransformer network. Our method was evaluated across various tasks, including indoor and outdoor semantic segmentation. Through extensive experiments on challenging benchmarks, thorough ablation studies and theoretical analysis, we demonstrated the robustness and effectiveness of our approach on real-world datasets. Our contributions significantly advance the state-of-the-art in-point cloud processing. The introduced techniques, including the Soft Masked Transformer, Skip Attention-based Up-sampling block, and Shared Point Position Encoding strategy, provide notable improvements in capturing intricate details and enhancing the performance of point cloud processing tasks. Experimental results on diverse datasets confirmed the efficacy of our proposed approach. As future work, exploring additional priors or refining the network architecture could offer promising avenues to further improve point cloud processing.
Limitations. Our current framework, like most point-based segmentation methods, operates under a closed-set assumption. When encountering novel classes or unseen geometric shapes, the absence of learned task-level priors may cause the model to misclassify them into the most semantically or geometrically similar known class. Addressing this limitation is crucial for real-world deployment. Future work may focus on integrating class-agnostic feature learning to capture generic geometric patterns, leveraging open-set recognition to detect unfamiliar categories during inference.